# UniDoc-Bench: A Unified Benchmark for Document-Centric Multimodal RAG

### v1

@@ 3.3 Dataset Quality
We evaluate whether our constructed dataset is of sufficient quality to support reliable evaluation of different RAG systems by sampling $20\%$ of our dataset— $40$ QA pairs from each domain, resulting in a total of $320$ QA pairs—for human evaluation. We recruited $3$ human annotators to evaluate the question–response pairs against the provided source documents. In cases where $3$ annotators disagreed, a $4$th senior reviewer mediated the discussion and guided the annotators toward a consensus decision. For each item, annotators were directed to a folder containing all relevant source materials, including text extracted from PDF documents and associated images. The annotation process involved assessing each question-response pair across five dimensions (More details about this human annotation task can be found in Appendix C):
@@
  • Grounding: assesses whether each source chunk (text, image, or table) used to generate the ground-truth response is necessary to answer the question, by labeling it as either required or not required.
@@
  • Human-like Intent: evaluates whether the question reflected a natural, meaningful query that a human would plausibly ask to retrieve information.
Table 2: Human evaluation quality on a $20\%$ sample (n=320n{=}320). Each cell shows % and (count/320)
Factuality–Q Factuality–R Completeness Self-Contained Human-like Intent Grounding
% & Count 99.70% (319/320) 91.90% (294/320) 91.90% (294/320) 99.70% (319/320) 97.50% (312/320) 84.38% (270/320)
As shown in Table 2, the sample shows near-perfect question factuality and self-containment, with strong response factuality and completeness (each $\approx$294/320). Human-like intent remains high (312/320). Grounding label accuracy is solid (270/320) as well.
@@ 4 Experiments
To fairly evaluate different RAG systems, we focus on two aspects: retrieval and end-to-end performance. In this section, we first evaluate the retrieval performance of four embedding and retrieval models, including text-only, image-only, and two multimodal approaches (Section 4.1). We then assess the end-to-end response performance of six RAG systems that vary in their use of embeddings, retrieval strategies, and LLMs (Section 4.2). Together, these experiments highlight the utility of our dataset and provide practical guidance for selecting RAG components.
@@ 4.1 Retrieval Performance
Baselines. We use the curated PDF documents as the knowledge base and the synthesized QA pairs to evaluate $4$ embedding–retrieval models. For all methods, we retrieve the top-$k=10$ candidates.
  • Text: PDF pages are parsed into text chunks, each embedded with OpenAI’s text-embedding-3-small, and retrieved via vector search.
  • Image: Each PDF page is converted to a JPEG image, which is embedded using ColQwen2.5-v0.2 (Faysse et al., 2024b) for image retrieval.
@@
    • MM (GME): Text and images are jointly embedded using GME-Qwen2-VL-7B -Instruct (Zhang et al., 2024a), enabling multimodal retrieval.
@@
Metrics. We report Precision@10 and Recall@10 as the retrieval metrics. Since no re-ranker is applied, recall is more informative than nDCG for evaluation. Since we need to evaluate both image and text retrieval, each retrieved text chunk or PDF image-page is mapped back to its original PDF page, and the ground-truth contexts are mapped in the same way. Consequently, a retrieved chunk may span multiple consecutive pages of the source document (e.g., pages 2–3 of document A). A retrieval is considered a true positive if the retrieved text chunk or image-page matches the ground-truth context in both page number and file. This criterion may slightly inflate Recall@10, since partial overlaps (e.g., retrieved pages 1–3 vs. ground-truth pages 3–5, with the answer on page 5) are still treated as correct. However, this approach offers the most practical and fair basis for comparing text and image retrieval. Thus, absolute scores should not be overinterpreted; the key is the relative performance differences across methods, which remain reliable.
Table 3: Retrieval performance (Precision@10 / Recall@10) of $4$ RAG systems on $1600$ QA pairs across eight domains, with average recall reported across all domains.
Domain Text (OpenAI) Image (colqwen) Multimodal
GME Text + Image
Precision Recall Precision Recall Precision Recall Precision Recall
Com. 0.430 0.813 0.294 0.831 0.354 0.895 0.523 0.886
Constr. 0.377 0.750 0.263 0.794 0.336 0.881 0.451 0.833
CRM 0.400 0.808 0.283 0.829 0.343 0.884 0.486 0.876
Edu 0.414 0.843 0.268 0.843 0.366 0.912 0.460 0.880
Energy 0.382 0.772 0.257 0.822 0.257 0.822 0.459 0.863
Fin. 0.384 0.778 0.291 0.812 0.376 0.857 0.484 0.867
HC 0.420 0.741 0.252 0.849 0.370 0.835 0.460 0.837
Legal 0.440 0.864 0.291 0.855 0.327 0.876 0.510 0.891
Avg. 0.406 0.796 0.275 0.829 0.341 0.870 0.479 0.867
Table 3 reports the retrieval performance of the four RAG embedding-retrieval models. We observe that image-based retrieval achieves consistently higher recall but lower precision than text-based retrieval, as page-image chunks cover more information than individual text chunks. Combining text and image retrieval (T++I) further improves both recall and precision, effectively leveraging the strengths of both modalities. In contrast, multimodal embeddings (gme-Qwen2-VL-7B-Instruct), which encode text and images jointly rather than separately, achieve comparable recall but substantially lower precision, suggesting that current multimodal embeddings still lag behind fusion of unimodal embeddings. We also break down retrieval performance by question and answer types in Appendix E.1.
@@ 4.2 End-to-End Performance
    • Image-only RAG (IMG): Uses LlamaIndex with colqwen2.5-v0.2 (Faysse et al., 2024b) for image retrieval and GPT-4.1 as the final MM-LLM. Each PDF page is converted to a JPEG image and embedded. After retrieval, the question and retrieved images are provided to GPT-4.1 to obtain the final response.
@@
    • VRAG (Wang et al., 2025d): a multimodal RAG agent that leverages a vision-specific action space—including operations such as cropping and scaling—to iteratively extract information from image-formatted PDF pages in a coarse-to-fine manner. The embedding model is colqwen2.5-v0.2, and the final LLM is GPT-4.1.
@@
  • Text-only RAG: Most multimodal RAG studies (Wang et al., 2025b; Faysse et al., 2024a) compare only against text-only baselines. For a fairer comparison, PDF pages are parsed into text chunks, embedded for retrieval, with associated images/tables linked back for final responses.
    • TEXT: Each text chunk is embedded using text-embedding-3-small and retrieved. The retrieved text chunks, along with their associated images, are then fed into GPT-4.1 to generate the final response.
    • Vertex AI: following the official tutorial [^4] https://www.cloudskillsboost.google/focuses/85643?parent=catalog, PDFs are parsed into text and images, with images auto-captioned by Gemini. Only the text (document text and image captions) is indexed by text-embedding-004 and retrieved, and the retrieved chunks along with the corresponding images are passed to gemini-2.5-flash for final response.
@@
  • MM-RAG: Both text chunks and image-format page images are embedded and retrieved for responses.
@@
    • Multimodal Text-Image-Fusion RAG (T++I): Retrieves text and images separately using text-embedding-3-small and colqwen2.5-v0.2, then combines them for generation with GPT-4.1.
@@
    • Multimodal-joint-Retrieval RAG (MM): Uses gme-Qwen2-VL-7B-Instruct (Zhang et al., 2024a) as a multimodal embedding model for both text and image content. Unlike T++I, where text and images are embedded and retrieved separately, the text chunks and image-formatted PDF pages are embedded together, retrieved jointly, and then fed into GPT-4.1 for the final response.
Table 4: Completeness of six RAG systems on 1,600 QA pairs across eight domains. Average recall is reported across all domains, with similarity top-$k$ set to 10 and 20, computed against the ground-truth responses.
Domain Image-only RAG Text-only RAG (+img matched) Multimodal RAG
IMG VRAG TEXT Vertex AI MM (GME) T+I
top-$10$ top-2020 top-$10$ top-2020 top-$10$ top-2020 top-$10$ top-2020 top-$10$ top-2020 top-$10$ top-2020
Com. 0.545 0.552 0.547 0.550 0.633 0.673 0.613 0.630 0.617 0.611 0.693 0.733
Constr. 0.502 0.601 0.536 0.542 0.561 0.587 0.558 0.621 0.616 0.609 0.607 0.647
CRM 0.524 0.524 0.523 0.544 0.643 0.663 0.628 0.625 0.623 0.637 0.647 0.703
Edu 0.569 0.560 0.517 0.524 0.692 0.702 0.613 0.633 0.640 0.668 0.688 0.691
Energy 0.535 0.566 0.558 0.589 0.607 0.637 0.627 0.677 0.669 0.666 0.649 0.680
Fin. 0.500 0.499 0.529 0.535 0.584 0.626 0.557 0.605 0.627 0.636 0.638 0.636
HC 0.481 0.492 0.481 0.492 0.602 0.639 0.638 0.643 0.642 0.664 0.621 0.666
Legal 0.558 0.568 0.599 0.595 0.629 0.696 0.642 0.675 0.609 0.629 0.689 0.716
Avg. 0.527 0.545 0.536 0.546 0.619 0.653 0.610 0.639 0.630 0.641 0.654 0.684
@@
Metrics. For end-to-end performance, we use an LLM-based judge to measure faithfulness and completeness. Specifically, we first ask the LLM to extract the facts required to answer each question and then verify whether these facts are grounded in the ground-truth chunks; this is measured as faithfulness. Next, we ask the LLM to extract the facts required to answer the question from the ground-truth answer and then check whether each fact appears in the system’s response; this is measured as completeness. Higher faithfulness and completeness scores are better.
@@
Table 4 reports the completeness of responses generated by the six RAG systems under varying similarity top-$k$ retrieval settings. Text-only RAG ($0.653$) substantially outperforms Image-only RAG systems (IMG: $0.545$, VRAG: $0.546$), highlighting the significant performance gap between text-based and image-based retrieval in current RAG architectures. Although image retrieval achieves higher completeness at the retrieval stage, this advantage does not translate into better end-to-end performance, since multimodal LLMs (GPT-4.1) are more effective when processing text and image chunks together rather than page-level image PDFs alone. The text-image-fusion RAG achieves the best overall performance ($0.684$) across eight domains, demonstrating that image-based PDF representations can effectively complement text retrieval. Although VRAG leverages cropping and scaling to enhance image-based retrieval ($0.536$ for VRAG vs. $0.527$ for IMG (topk=10k=10)), it still lags behind the combined Text&Image-Retrieval approach, underscoring the advantage of explicitly integrating both modalities. Multimodal joint-retrieval RAG systems (MM; $0.641$) also fall short of the simple combination of the best text and image embeddings. This indicates that current multimodal embedding approaches still have substantial room for improvement, and that explicitly combining separate text and image embeddings remains the most effective strategy for leveraging multimodal documents. More notably, multimodal-joint RAG (MM; $0.641$) performs worse than text-only RAG ($0.653$), demonstrating that current multimodal models still fall short of strong unimodal baselines. These results also highlight the importance of establishing fair baselines and the value of our dataset: multimodal RAG systems should be benchmarked against strong, balanced baselines on diverse and high-quality datasets rather than against overly weak text-only settings.
@@
Table 5 shows that questions requiring only text are most effectively handled by RAG systems with text-embedding. Questions requiring tables are also relatively easy for RAG systems, as tables can be accurately parsed as text, which is a straightforward step before embedding documents for text-based retrieval. In contrast, questions requiring images remain challenging across all embedding types — text, image, or multimodal — highlighting that future RAG improvements should prioritize image-required questions. We also observe that question type has minimal impact on overall RAG performance. We provide detailed case studies in Appendix D.
Table 5: Faithfulness and Completeness of six RAG systems across different question and answer types on 1,600 QA pairs spanning eight domains, with average recall reported across all domains.
Type Image-only RAG Text-only RAG (+img matched) Multimodal RAG
IMG VRAG TEXT Vertex AI MM (GME) T+I
faith. complet. faith. complet. faith. complet. faith. complet. faith. complet. faith. complet.
F.R. 0.640 0.536 0.581 0.536 0.698 0.629 0.563 0.557 0.668 0.599 0.763 0.704
Comp. 0.669 0.510 0.611 0.513 0.739 0.619 0.634 0.644 0.744 0.656 0.755 0.641
Summary 0.727 0.536 0.706 0.602 0.736 0.613 0.694 0.670 0.752 0.670 0.781 0.651
Logical 0.738 0.526 0.650 0.584 0.769 0.607 0.690 0.660 0.744 0.678 0.780 0.621
Text-only 0.812 0.580 0.767 0.624 0.877 0.656 0.817 0.758 0.849 0.771 0.880 0.700
Img-only 0.512 0.448 0.453 0.483 0.580 0.606 0.359 0.447 0.463 0.436 0.620 0.615
Text + Img 0.678 0.498 0.576 0.523 0.716 0.601 0.581 0.556 0.707 0.583 0.749 0.630
Table-req. 0.693 0.587 0.662 0.554 0.714 0.601 0.716 0.670 0.819 0.747 0.811 0.716
4.3 Additional findings
MM-RAG systems can offer both improved end-to-end performance and lower cost compared to text-only RAG. As reported in Appendix E.3, text-only RAG is the most expensive, image-only RAG has the lowest cost and latency, and multimodal RAG is cheaper than text-only RAG while maintaining comparable latency.
Open-source and commercial multimodal embeddings perform comparably. We compare RAG systems using different multimodal embeddings (Table 8, Table 9) and find that the commercial voyage-multimodal-3 achieves similar performance to the open-source GME, though both still lag behind multimodal text–image fusion RAG systems.
Content-rich images increase difficulty. We classify images using gemini-2.5-pro as content-rich (containing information not in the text) or illustrative. Content-rich images are more prevalent in finance (62.8%) and construction (69.3%) than in commerce manufacturing (40.0%) and legal (49.5%), indicating that domains with more content-rich images pose greater challenges for RAG, consistent with the results in Table 4. Details are in Appendix F.1 .
Question type affects difficulty. We further analyzed fined-grained evidence types and found that RAG performance depends on answer modality: text retrieval excels at entity recognition (53.9% better than image retrieval), comparative analysis (37.6%), contextual numerical reasoning (34.8%), and quantity estimation (29.1%), while image retrieval is stronger on chart/table interpretation (64.2% better than text retrieval), temporal trends (40.0%), and spatial/geographic reasoning (13.3%). Detailed examples and analysis are in Appendix D.1, Appendix D.2 and Appendix F.2.
We also summarize in Appendix F that single document page numbers and formats do not significantly affect MM-RAG performance.

### v3

@@ 3.3 Dataset Quality
We evaluate whether our UniDoc-Bench is of sufficient quality to support reliable evaluation of different RAG systems by recruiting $5$ human annotators to evaluate the $1,600$ question–response pairs against the provided source documents. The annotation process involved assessing each question-response pair across five dimensions (Appendix C):
@@
  • Grounding: assesses whether each source chunk (text, image, or table) used to generate the ground-truth response is required to answer the question, by labeling it as either required or not required, and these labels serve as the ground truth. We then compare the labels produced by our pipeline against the human-annotated ground truth to compute accuracy.
@@
  • Human-like Intent: evaluates whether the question reflected a natural, meaningful query that a human would ask to retrieve information.
Table 2: Human evaluation quality on the 1,6001{,}600 QAs.
Fact.–Q Fact.–R Complete.
(%) 98.61 94.20 93.63
Self-Cont. Human-like Grounding
(%) 98.25 96.25 84.38
As shown in Table 2, the sample shows near-perfect question factuality and self-containment, with strong response factuality and completeness. Human-like intent remains very high ($96.25\%$). Grounding label accuracy is also solid ($84.38\%$). Any questions or responses that do not receive uniformly positive labels are revised by human annotators. These results demonstrate the high quality of UniDoc-Bench for evaluating MM-RAG systems, as well as the robustness of our synthesis pipeline, which can be readily used to generate reliable QA pairs for new databases.
Dataset Statistics. UniDoc-Bench consists of $200$ QA pairs for each domain, in total $1600$ human-verified and revised QAs. Within each domain, we have an equal distribution of $50$ text-only, image-only, text-plus-image, and table-required questions. In total, the dataset contains $800$ single-modality and $800$ multi-modality questions. On average, each question requires $2.15$ evidence items (text chunks, images, or tables) for a complete answer. More details can be found in Figure 1(b).
@@ 4 Experiments
To fairly evaluate different RAG systems, we focus on two aspects: retrieval and end-to-end performance. In this section, we first evaluate the retrieval performance of $4$ embedding and retrieval models, including text-only, image-only, and two multimodal approaches (§ 4.1). Next, we evaluate the end-to-end response quality of nine RAG systems that differ in their embeddings, retrieval strategies, and underlying LLMs (Section 4.2). Finally, we demonstrate how our dataset can be used for VQA tasks (Section 4.3). Together, these experiments highlight the usefulness of our dataset and provide practical guidance for selecting RAG components and evaluating VQA systems.
@@ 4.1 Retrieval Performance
Baselines. We use the curated PDF documents as the knowledge base and the synthesized $1,600$ QA pairs to evaluate the following $4$ embedding–retrieval models. For all methods, we retrieve the top-$k=10$ candidates.
  • Text: PDFs are parsed11footnotemark: 1 into text chunks, each embedded with OpenAI’s text-embed ding-3-small, and retrieved via vector search.
  • Image: Each PDF page is converted to an image, which is embedded using ColQwen 2.5-v0.2 (faysse2024colpaliefficientdocumentretrieval) for image retrieval.
@@
    • MM (GME): Text and images are jointly embedded using GME-Qwen2-VL-7B -Instruct (gme), enabling multimodal retrieval.
@@
Metrics. We report Precision@10 and Recall @10 as the retrieval metrics. Since no re-ranker is applied, recall is more informative than nDCG for evaluation. Since we need to evaluate both image and text retrieval, each retrieved text chunk or PDF image-page is mapped back to its original PDF page, and the ground-truth contexts are mapped in the same way. Consequently, a retrieved chunk may span multiple consecutive pages of the source document (e.g., pages 2–3 of document A). A retrieval is considered a true positive if the retrieved text chunk or image-page matches the ground-truth context in both page number and file. This criterion may slightly inflate Recall@10, since partial overlaps (e.g., retrieved pages 1–3 vs. ground-truth pages 3–5, with the answer on page 5) are still treated as correct. However, this approach offers the most practical and fair basis for comparing text and image retrieval. Thus, absolute scores should not be overinterpreted; the key is the relative performance differences across methods.
Table 3: Retrieval performance (Precision@10 / Recall@10) of four RAG systems on 1,600 QA pairs across eight domains (top) and broken down by question and answer types (bottom).
Domain Text Image MM (GME) T++I
Prec. Rec. Prec. Rec. Prec. Rec. Prec. Rec.
Com. .286 .829 .179 .843 .437 .882 .449 .914
Cons. .246 .762 .159 .792 .429 .864 .422 .853
CRM .271 .783 .175 .830 .437 .860 .426 .863
Edu .278 .855 .160 .851 .432 .878 .427 .896
Energy .239 .706 .148 .718 .366 .723 .374 .746
Fin. .254 .781 .177 .818 .434 .891 .427 .912
HC .297 .746 .151 .856 .455 .859 .455 .851
Legal .312 .861 .178 .861 .462 .883 .458 .903
Avg. .273 .790 .166 .821 .431 .855 .430 .864
By Question Type
F.R. .205 .747 .140 .825 .226 .859 .219 .869
Comp. .283 .820 .163 .835 .313 .901 .309 .909
Summary .336 .828 .200 .800 .355 .880 .360 .895
Logical .365 .820 .201 .813 .386 .870 .382 .882
By Answer Type
Text-only .383 .839 .217 .790 .406 .847 .400 .849
Img-only .081 .724 .092 .878 .097 .909 .097 .920
Text + Img .336 .847 .190 .824 .350 .888 .351 .908
Table-req. .291 .752 .163 .791 .326 .851 .316 .861
Table 3 summarizes the retrieval performance of four RAG embedding–retrieval models across eight domains, four question types, and four answer types We observe that image-based retrieval achieves consistently higher recall but lower precision than text-based retrieval, as page-image chunks cover more information than individual text chunks. Combining text and image retrieval (T++I) further improves both recall and precision, effectively leveraging the strengths of both modalities. In contrast, multimodal embeddings (GME-Qwen2-VL-7B-Instruct), which encode text and images jointly rather than separately, achieve comparable precision but lower recall, suggesting that current multimodal embeddings still lag behind fusion of unimodal embeddings.
@@ 4.2 End-to-End Performance
    • Image-only RAG (IMG): Uses LlamaIndex with colqwen2.5-v0.2 (faysse2024colpaliefficientdocumentretrieval) for image retrieval. After retrieval, the question and retrieved images are provided to GPT-4.1 to obtain the final response.
@@
    • VRAG (wang2025vragrlempowervisionperceptionbasedrag): a multimodal RAG agent that uses a vision-specific action space — cropping and scaling — to iteratively extract information from image-formatted PDF pages in a coarse-to-fine manner. The embedding model is colqwen2.5-v0.2, and the final LLM is GPT-4.1.
@@
  • Text-only RAG (TEXT): Most multimodal RAG studies (wang2025vidorag; faysse2024colpali) compare only against text-only baselines. For a fairer comparison, PDF pages are parsed into text chunks, embedded for retrieval, with associated images/tables linked back for final responses. In this baseline, each text chunk is embedded using text-embedding-3-small and retrieved. The retrieved text chunks, along with their associated images, are then fed into GPT-4.1 to generate the final response.
@@
  • MM-RAG: Both parsed text and image-format page images are embedded and retrieved.
@@
    • Multimodal Text-Image-Fusion RAG (T++I): Retrieves text and images separately using text-embedding-3-small and colqwen2.5-v0.2, then combines them for generation with GPT-4.1. We also evaluate multiple state-of-the-art LLMs, including Gemini-pro-2.5, Claude 4.5, and GPT-5.
@@
    • Multimodal-joint-Retrieval RAG (MM): Uses GME-Qwen2-VL-7B-Instruct (gme) (MM(G)) or voyage -multimodal-3 (MM(V)) as a multimodal embedding model for both text and images.Unlike T++I, where text and images are embedded and retrieved separately, the text chunks and image-formatted PDF pages are embedded together, retrieved jointly, and then fed into GPT-4.1 for the final response.
Table 4: Completeness of systems evaluated on 1,600 QA pairs across $8$ domains. Average recall is reported over all domains, with similarity top-$k$ set to 10. Gemini refers to Gemini-2.5-pro. Claude refers to Claude-4.5-sonnet. For VQA, the first value uses the entire document as image input, while the second value uses ground-truth images only. GT is the performance of Claude-4.5-sonnet on the ground-truth text chunks, images and tables.
Domain Image-only Text-only Multimodal VQA GT
IMG VRAG TEXT MM (V) MM (G) T+I
GPT-4.1 GPT-4.1 Gemini Claude GPT-5 Gemini Claude GPT-5 Claude
Com. .545 .547 .633 .663 .657 .693 .707 .789 .746 .613/.665 .670/.805 .629/.706 .883
Cons. .502 .536 .561 .600 .592 .607 .662 .737 .648 .566/.610 .669/.706 .597/.627 .776
CRM .524 .523 .643 .635 .640 .647 .689 .771 .696 .612/.679 .756/.774 .614/.666 .848
Edu .569 .517 .692 .660 .673 .688 .672 .765 .637 .612/.636 .720/.741 .612/.651 .845
Energy .535 .558 .607 .675 .661 .649 .682 .768 .721 .584/.710 .750/.799 .646/.679 .830
Fin. .500 .529 .584 .641 .638 .638 .672 .788 .693 .585/.635 .727/.808 .631/.670 .835
HC .481 .481 .602 .628 .651 .621 .689 .767 .665 .580/.673 .723/.735 .604/.647 .849
Legal .558 .599 .629 .597 .600 .689 .705 .770 .714 .636/.654 .647/.740 .671/.680 .858
Avg. .527 .536 .619 .637 .639 .654 .685 .770 .690 .599/.658 .708/.763 .625/.666 .840
By Question Type
F.R. .557 .344 .648 .619 .612 .677 .687 .739 .694 .569/.685 .709/.756 .645/.698 .829
Comp. .542 .418 .633 .638 .646 .641 .700 .792 .683 .516/.660 .722/.768 .638/.662 .825
Summary .536 .407 .626 .652 .649 .640 .666 .759 .689 .530/.638 .670/.777 .596/.627 .867
Logical .548 .513 .637 .664 .681 .630 .681 .813 .706 .514/.602 .719/.774 .607/.639 .864
By Answer Type
Text-only .588 .464 .700 .777 .771 .695 .767 .863 .773 .582/.680 .752/.823 .676/.691 .923
Img-only .486 .336 .616 .465 .462 .619 .588 .644 .629 .510/.613 .577/.651 .546/.611 .756
Text+Img .502 .453 .600 .584 .580 .617 .611 .719 .617 .441/.587 .677/.729 .578/.609 .813
Table-req. .610 .392 .633 .723 .742 .683 .773 .853 .741 .609/.752 .812/.851 .700/.752 .870
@@
Metrics. For end-to-end performance, we use an LLM-based judge to measure faithfulness and completeness. Specifically, we first ask the LLM to extract the facts required to answer each question and then verify whether these facts are grounded in the ground-truth chunks; this is measured as faithfulness ($\uparrow$). Next, we ask the LLM to extract the facts required to answer the question from the ground-truth answer and then check whether each fact appears in the system’s response; this is measured as completeness ($\uparrow$).
@@
Table 4 (red background) reports the completeness of responses generated by the six RAG systems. Text-only RAG ($0.619$) substantially outperforms Image-only RAG systems (IMG: $0.527$, VRAG: $0.536$), highlighting the significant performance gap between text-based and image-based retrieval in current RAG architectures. Although image retrieval achieves higher recall at the retrieval stage, this advantage does not translate into better end-to-end performance, since multimodal LLMs (GPT-4.1) are more effective when processing text and image chunks together rather than page-level image PDFs alone. In addition, the low precision of image retrieval makes it harder for the model to identify the correct information. The text-image-fusion RAG (T++I) achieves the best overall performance ($0.654$) across eight domains, demonstrating that image-based PDF representations can effectively complement text retrieval. Although VRAG leverages cropping and scaling to enhance image-based retrieval ($0.536$ for VRAG vs.0.5270.527 for IMG), it still lags behind the combined T++I approach, underscoring the advantage of explicitly integrating both modalities. Multimodal joint-retrieval RAG systems (MM (voyage-multimodal-3): $0.637$; MM (GME-Qwen2-VL-7B-Instruct): $0.639$) also fall short of the simple combination of the best text and image embeddings. This indicates that current multimodal embedding approaches still have substantial room for improvement, and that explicitly combining separate text and image embeddings remains the most effective strategy for leveraging multimodal documents. More notably, in some domains—CRM, Education and Legal—multimodal joint RAG performs worse than text-only RAG, indicating that current multimodal models still lag behind strong unimodal baselines in certain domains. These results highlight the importance of establishing fair baselines and the value of UniDoc-Bench: multimodal RAG systems should be benchmarked against strong, balanced baselines on diverse and high-quality datasets rather than against overly weak text-only settings.
@@
Table 4 (column T+I) compares different state-of-the-art LLMs used in the Text&Image Retrieval setting. Claude-4.5-sonnet achieves the best performance across all domains, question types, and answer types. The table also shows that questions requiring only text are most effectively handled by RAG systems with text-embedding. Questions requiring tables are also relatively easy for RAG systems, as tables can be accurately parsed as text, which is a straightforward step before embedding documents for text-based retrieval. In contrast, questions requiring images remain challenging across all embedding types — text, image, or multimodal — highlighting that future RAG improvements should prioritize image-required questions. We further observe that multimodal joint RAG achieves stronger performance on text-dominant questions, whereas the T++I RAG is more effective for image-dominant queries. We also provide detailed case studies in Appendix D.
4.3 Visual Question Answering Performance
UniDoc-Bench can also be used to evaluate Visual Question Answering (VQA) tasks. Table 4 (gray background) reports the performance of state-of-the-art LLMs — Gemini-pro-2.5, Claude -4.5-Sonnet, and GPT-5 — when applied to entire image-format PDFs and to ground-truth pages only. The results show that Claude-4.5-Sonnet consistently achieves the highest completeness scores across all domains and question types in the VQA setting. All models exhibit a performance gap between the two settings, confirming that reasoning over entire documents is more challenging than over isolated ground-truth images. Gemini-pro-2.5 is the most sensitive to this noise. In contrast, Claude-4.5-Sonnet and GPT-5 are more robust to full-document inputs, showing smaller performance drops.
Additional Findings. We show the best performance of Claude-4.5-sonnet on the ground-truth chunks in “GT” column of Table 4. Cost and latency are reported in Appendix E.1. Case studies on the impact of content-rich images are presented in Appendix F.1. Analyses of how question type affects difficulty are provided in Appendix D.1, D.2, and F.2. Finally, Appendix F shows that the number of pages and document formats do not significantly affect MM-RAG performance.